Dipole is placed parallel to the electric field. If w is the work done in rotating the dipole by 60 °, then work done in rotating it by 180° is.

Grade:12

2 Answers

Divya Chauhan
6 years ago
w=-PE(cos60°-cos0°) so w = PE/2 ; now work done in rotating the dipole by 180°(w1)=-PE(cos180°-cos 60°); w1=3*PE/2; so w1=3w.
Mehak brar
5 years ago
The point charges Q and _  2Q are placed some distance   apart if the electric field at the location of Q is E , then find d the electric at the location of _2Q